# Certifiable Deep Learning for Reachability Using a New Lipschitz Continuous Value Function

### v2

@@ V Learning the New RA Value Function
Motivated by the fact that the optimal RA control policy is deterministic [36], we adapt max-min Deep Deterministic Policy Gradient (DDPG) [42], a deep RL method for learning deterministic policies and their value functions, to learn ${\pi}$, ${\phi}$ and ${V_{\gamma}(x)}$.
Let $\gamma$ be an arbitrary time discount factor in $(0,1)$. Similar to prior works [11, 15, 12], we approximate the optimal control policy $u={\pi}^{*}(x)$ and the worst-case disturbance policy $d={\phi}^{*}(x,{\pi}^{*}(x))$ by neural network (NN) policies $u={\pi_{\theta_{u}}}(x)$ and $d={\phi_{\theta_{d}}}(x)$, respectively, with ${\theta_{u}}$ and ${\theta_{d}}$ being their parameters. We define an NN Q function as ${Q_{\theta_{q}}}(x,u,d):\mathbb{R}^{n}\times{\mathcal{U}}\times{\mathcal{D}}\to\mathbb{R}$, where ${\theta_{q}}$ represents the NN’s parameter vector. Substituting ${\pi_{\theta_{u}}}$ and ${\phi_{\theta_{d}}}$ into ${Q_{\theta_{q}}}$, we can derive an NN value function ${V_{\theta}}(x):={Q_{\theta_{q}}}(x,{\pi_{\theta_{u}}}(x),{\phi_{\theta_{d}}}(x))$, where $\theta$ is the concatenation of parameters ${\theta_{q}}$, ${\theta_{u}}$ and ${\theta_{d}}$. Let $\mathbb{P}$ be the sampling distribution in learning. In max-min DDPG, we learn ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ by jointly optimizing the following problems:
@@ VI-A Certification using Lipschitz constants
and a state trajectory under ${\bar{u}}_{t}$ using an arbitrary $d_{t}\in{\mathcal{D}}$:
@@
Note that if we can verify that a trajectory starting at state $x$ reaches the target set safely despite any disturbance within $T$ stages, then it suffices to claim $x\in{\mathcal{R}}$. This certification horizon $T$ can be set arbitrarily. Ideally, we would set $T=\infty$, but it is impractical to evaluate an infinitely long trajectory. Therefore, during certification, we consider a finite, user-defined $T$. This $T$ should preferably be long enough to allow initial states to reach the target set. A short $T$ results in a conservative certification since it overlooks the possibility that the trajectory might safely reach the target set at a later time.
We assume that the dynamics $f(x,u,d)$ is Lipschitz continuous and there exists an upper bound on the disturbance in ${\mathcal{D}}$ at each stage $t$, i.e., $\|d_{t}\|_{2}\leq{\varepsilon_{d}}$. Let ${L_{f_{x}}}$ and ${L_{f_{d}}}$ be the Lipschitz constants of the dynamics $f$ with respect to the state $x$ and disturbance $d$, respectively. At time $t=1$, we observe
$\|{\bar{x}}_{1}$ $-{\tilde{x}}_{1}\|_{2}\leq\|f({\bar{x}}_{0},{\bar{u}}_{0},0)-f({\tilde{x}}_{0},{\bar{u}}_{0},d_{0})\|_{2}$
@@
Thus, when ${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)>0$, the set ${\mathcal{E}}_{{\bar{x}}_{0}}$ is certified to be within the ground truth RA set ${\mathcal{R}}$. Moreover, $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs that can drive all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite any disturbance in ${\mathcal{D}}$.
@@ VI-B Certification using second-order cone programming
First, we formulate a surrogate RA problem: A subset of the original target set ${\mathcal{T}}$, represented by the interior of a polytope ${\check{\mathcal{T}}}:=\{x:P_{i}x-k_{i}>0,i\in\mathcal{I}_{{\check{\mathcal{T}}}}\}$, is defined as a surrogate target set ${\check{\mathcal{T}}}$, where $\mathcal{I}_{\check{\mathcal{T}}}$ is a finite index set of the polytope’s edges. Additionally, we define a subset of the original constraint set ${\mathcal{C}}$, represented by the intersection of a finite number of quadratic functions’ super-zero level sets ${\check{\mathcal{C}}}:=\{x:\frac{1}{2}x^{\top}Q_{i}x+q_{i}^{\top}x+b_{i}>0,Q_{i}\succeq 0,i\in\mathcal{I}_{{\check{\mathcal{C}}}}\}$, as a surrogate (nonconvex) constraint set ${\check{\mathcal{C}}}$. The set ${\check{\mathcal{C}}}$ could be nonconvex when approximating collision avoidance constraints.
Subsequently, we leverage SOCPs to verify if a state trajectory from $x_{0}$ can reach ${\check{\mathcal{T}}}$ while staying within ${\check{\mathcal{C}}}$ despite any disturbance. We achieve this by sequentially minimizing each function defined in ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$, iterating from the stage $\tau=0$ to $\tau=T$. If their minimum is positive at a stage $t$ and there is no intermediate stage $\tau<t$ such that $x_{\tau}$ is outside ${\check{\mathcal{C}}}$, we claim that the initial state $x_{0}$ can safely reach the original target set ${\mathcal{T}}$, under all possible disturbances.
@@
where the bounds on the dynamics $f$ can be derived using a Taylor series. At a stage $t$, we can verify if $x_{t}\in{\check{\mathcal{T}}}$ by solving a sequence of SOCPs iterating over all $i\in\mathcal{I}_{\check{\mathcal{T}}}$, and checking if their minimum is positive:
${\check{r}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{T}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}$ $P_{i}^{\top}x_{t}-k_{i}\big{\}}.$ (23)
@@
${\check{c}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{C}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\frac{1}{2}x_{t}^{\top}Q_{i}x_{t}+q_{i}^{\top}x_{t}+b_{i}\big{\}}.$ (24)
@@
and $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs, capable of driving all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite any disturbance.
Running example (continued). We evaluate ${\check{r}_{t}^{S}}=\min_{i}\{\check{r}_{t,i}^{S}\}_{i=1}^{6}$ by solving the following SOCPs, where each $\check{r}_{t,i}^{S}$ corresponds a function in the definition of ${\mathcal{T}}$ in (5):
@@
$\check{r}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3-p_{y,t}^{1},$ $\check{r}_{t,6}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3+p_{y,t}^{1}$
@@
$\check{c}_{t,i}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}(-1)^{i}\times p_{z,t}^{1}-p_{y,t}^{1}+0.05,\hskip 5.69046pti\in\{3,4\},$
We overapproximate the maximum height difference between two drones at the stage $t$ via $\Delta_{z,t}^{21}:=\max_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}p_{z,t}^{2}-p_{z,t}^{1}$, and consider
$\check{c}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\mathopen{}\mathclose{{}\left\|\begin{bmatrix}p_{x,t}^{1}-p_{x,t}^{2}\\ p_{y,t}^{1}-p_{y,t}^{2}\end{bmatrix}}\right\|_{2}^{2}-(1+\max(\Delta_{z,t}^{21},0))\times 0.2$ (27)
Remark 1 (Comparing the two methods).
1) SOCP certification can be less conservative than Lipschitz certification when ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$ can represent ${\mathcal{T}}$ and ${\mathcal{C}}$ exactly, and ${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}$ is a subset of the Lipschitz dynamically feasible set ${\mathcal{X}_{t,\bar{x}_{0}}^{L}}$, as defined in Section VI-A, for all $t\leq T$. This is because Lipschitz certification adds extra conservatism when estimating lower bounds of $r(x)$ and $c(x)$ using the Lipschitz constant, as shown in (19) and (20); 2) However, Lipschitz certification is faster to compute than SOCP because calculating ${\check{r}_{t}^{L}}$ and ${\check{c}_{t}^{L}}$ is easier than evaluating ${\check{r}_{t}^{S}}$ and ${\check{c}_{t}^{S}}$.
@@
VI-C Combining reachability learning and certification for computing trustworthy RA sets
@@
Online certification: Let $x$ be an arbitrary state. We can use the RA certificates ${\check{V}^{L}_{\gamma}}(x,T)$ in (21) or ${\check{V}^{S}_{\gamma}}(x,T)$ in (25) as online RA certification methods, verifying if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq{\varepsilon_{x}}\}$ can reach the target set safely despite any disturbance. We can compute ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ in real-time (10 Hz or faster), as shown in Figure 5. It can also be integrated into the safety filter proposed in [11], and our method provides a more robust RA certification than [11] as we certify whether all elements in ${\mathcal{E}}_{x}$ can reach the target set safely, allowing us to verify the RA capability without perfect state estimation.
Offline certification: We consider a finite set of states :=$\mathcal{L}:=\{x^{(i)}\}_{i=1}^{N}$ such that the union of their neighboring sets ${\mathcal{E}}_{x^{(i)}}=\{x:\|x-x^{(i)}\|_{2}\leq{\varepsilon_{x}}\}$ covers the set of states that we aim at certifying. For example, this includes the area near the orange gate in drone racing, as show in Figure 1. We enumerate each element $x\in\mathcal{L}$ and certify whether $\mathcal{E}_{x}\subseteq{\mathcal{R}}$ by checking if ${\check{V}^{L}_{\gamma}}$ in (21) or ${\check{V}^{S}_{\gamma}}$ in (25) is positive. The union ${\mathcal{S}}$ of those certified sets constitutes a subset of the ground truth RA set ${\mathcal{R}}$. For all elements in ${\mathcal{S}}$, we guarantee that they can reach the target set safely amid potential disturbances.
@@
Algorithm 1 Reachability Learning and Certifications:
VII Experiments
@@
We test our reachability learning and certification methods [^1] Experiment code and details are available at https://github.com/jamesjingqili
/Lipschitz_Continuous_Reachability_Learning.git
@@
in a 12-dimensional drone racing hardware experiment, and a triple-vehicle highway take-over simulation. In the highway simulation, we control one ego vehicle, modeled with nonlinear unicycle dynamics, to safely overtake another vehicle while avoiding a third vehicle driving in the opposite direction, as shown in Figure 4. Figures 1 and 4 demonstrate the high quality of the learned and certified RA sets. Additionally, the hardware experiments in Figure 3 show that our learned control policy ${\pi_{\theta_{u}}}$ safely controls the ego drone (highlighted with brightness) to overtake the other drone when the initial states of both agents are within the SOCP certified set.
@@
VII-A Hypothesis 1: Our learned policy has a higher success rate than state-of-the-art constrained RL methods
@@
We compare our learned policy ${\pi_{\theta_{u}}}$ with Deep Deterministic Policy Gradient-Lagrangian (DDPG-L)[34], Soft Actor Critic-Lagrangian (SAC-L) [35], Constrained Policy Optimization (CPO) [33]. We summarize the results in Table I. The success rate is estimated by computing the ratio of sampled initial states that can reach the target set safely under randomly generated disturbances. Our method achieves a 1.0 success rate when the initial states are sampled from the SOCP certified set, validating the deterministic guarantee that all elements in the certified sets can safely reach the target set despite any disturbance. However, CPO fails to converge for the drone racing experiment due to the complex and nonconvex constraints.
@@
TABLE I: Success rates table. Our method achieves a $1.0$ success rate when the initial states are sampled from the SOCP certified set.
Figure 3: In both subfigures, the initial states are within the SOCP certified set shown in Figure 1, validating the deterministic guarantees that the ego drone can safely overtake the other despite any disturbance on the other drone’s acceleration, with $\|d_{t}\|_{2}\leq 0.1m/s^{2}$. The remaining 9-dimensional initial state includes $[v_{x,t}^{1},v_{y,t}^{1},v_{z,t}^{1},p_{x,t}^{2},p_{y,t}^{2},p_{z,t}^{2},v_{x,t}^{2},v_{y,t}^{2},v_{z,t}^{2}]=[0,0.7,0,0.4,-2.2,0,0,0.3,0]$.
@@
Figure 4: Highway reachability analysis: In (a), we simulate the nonlinear dynamics with the learned policy ${\pi_{\theta_{u}}}$ and randomly sampled disturbances on other vehicles’ acceleration. The 10-dimensional state space includes $[p_{x,t}^{1},p_{y,t}^{1},v_{t}^{1},\theta_{t}^{1},p_{x,t}^{2},p_{y,t}^{2},v_{y,t}^{2},p_{x,t}^{3},p_{y,t}^{3},v_{y,t}^{3}]$. In (b), we project our learned value function, with $\gamma=0.95$, onto the $(x,y)$ position of the ego vehicle. In (c), we plot the RA set learned using the state-of-the-art method [12, 15] with $\gamma=0.95$. As suggested in [8], annealing $\gamma\to 1$ is necessary for prior works; otherwise, the learned RA sets in prior works are conservative. In (d), we plot our certified RA sets.
@@
VII-B Hypothesis 2: Our online RA set certification methods can be computed in real-time
Figure 5 shows that ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ can be computed in real-time to certify if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq 0.1\}$ can safely reach the target set, amid all potential disturbances. This enables real-time online certification.

### v5

@@ V Learning the New RA Value Function
Since the optimal RA control policy is deterministic [37], we adapt max-min Deep Deterministic Policy Gradient (DDPG) [43], a deep RL method for learning deterministic policies and their value functions, to learn ${\pi}$, ${\phi}$ and ${V_{\gamma}}$.
Let $\gamma$ be an arbitrary time discount factor in $(0,1)$. Similar to prior works [10, 14, 11], we approximate the optimal control policy ${\pi}^{*}(x)$ and the worst-case disturbance policy ${\phi}^{*}(x,{\pi}^{*}(x))$ by neural network (NN) policies ${\pi_{\theta_{u}}}(x)$ and ${\phi_{\theta_{d}}}(x)$, respectively, with ${\theta_{u}}$ and ${\theta_{d}}$ being their parameters. We define an NN Q function as ${Q_{\theta_{q}}}:\mathbb{R}^{n}\times{\mathcal{U}}\times{\mathcal{D}}\to\mathbb{R}$, where ${\theta_{q}}$ represents the NN’s parameter vector. Substituting ${\pi_{\theta_{u}}}$ and ${\phi_{\theta_{d}}}$ into ${Q_{\theta_{q}}}$, we can derive an NN value function ${V_{\theta}}(x):={Q_{\theta_{q}}}(x,{\pi_{\theta_{u}}}(x),{\phi_{\theta_{d}}}(x))$, where $\theta$ is the concatenation of parameters ${\theta_{q}}$, ${\theta_{u}}$ and ${\theta_{d}}$. Let $\mathbb{P}$ be a sampling distribution with a sufficiently large support in $\mathbb{R}^{n}$ that covers at least a part of the target set. In max-min DDPG, we learn ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ by alternatively optimizing the following problems:
@@ VI-A Certification using Lipschitz constants
and a disturbed state trajectory $\{{\tilde{x}}_{t}\}_{t=0}^{T}$ under $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ using an arbitrary $d_{t}\in{\mathcal{D}}$, $\forall t=0,\dots,T-1$:
@@
Note that if we can verify that a trajectory starting at state $x$ reaches the target set safely despite disturbances within $T$ stages, then it suffices to claim $x\in{\mathcal{R}}$, where the certification horizon $T$ can be set arbitrarily. Ideally, we would set $T=\infty$, but it is impractical to evaluate an infinitely long trajectory. Therefore, during certification, we consider a finite, user-defined $T$. This $T$ should preferably be long enough to allow initial states to reach the target set. A short $T$ results in a conservative certification since it overlooks the possibility that the trajectory might safely reach the target set at a later time.
We assume that the dynamics $f$ is Lipschitz continuous and there exists an upper bound on the disturbance in ${\mathcal{D}}$ at each stage $t$, i.e., $\|d_{t}\|_{2}\leq{\varepsilon_{d}}$. Let ${L_{f_{x}}}$ and ${L_{f_{d}}}$ be the Lipschitz constants of the dynamics $f$ with respect to the state $x$ and disturbance $d$, respectively. At time $t=1$, we observe
$\|{\bar{x}}_{1}$ $-{\tilde{x}}_{1}\|_{2}{=}\|f({\bar{x}}_{0},{\bar{u}}_{0},0)-f({\tilde{x}}_{0},{\bar{u}}_{0},d_{0})\|_{2}$
@@
Thus, when ${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)>0$, the set ${\mathcal{E}}_{{\bar{x}}_{0}}$ is certified to be within the ground truth RA set ${\mathcal{R}}$. Moreover, $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs that can drive all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite disturbances in ${\mathcal{D}}$.
@@ VI-B Certification using second-order cone programming
First, we formulate a surrogate RA problem: A subset of the original target set ${\mathcal{T}}$, represented by the interior of a polytope ${\check{\mathcal{T}}}:=\{x:P_{i}x-k_{i}>0,i\in\mathcal{I}_{{\check{\mathcal{T}}}}\}$, is defined as a surrogate target set ${\check{\mathcal{T}}}$, where $\mathcal{I}_{\check{\mathcal{T}}}$ is a finite index set of the polytope’s edges. Additionally, we define a subset of the original constraint set ${\mathcal{C}}$, represented by the intersection of a finite number of positive semidefinite quadratic functions’ super-zero level sets ${\check{\mathcal{C}}}:=\{x:\frac{1}{2}x^{\top}Q_{i}x+q_{i}^{\top}x+b_{i}>0,i\in\mathcal{I}_{{\check{\mathcal{C}}}}\}$, as a surrogate (nonconvex) constraint set ${\check{\mathcal{C}}}$. The set ${\check{\mathcal{C}}}$ could be nonconvex when approximating collision avoidance constraints.
Subsequently, we leverage SOCPs to verify if a state trajectory from $x_{0}$ can reach ${\check{\mathcal{T}}}$ while staying within ${\check{\mathcal{C}}}$ despite disturbances. We achieve this by sequentially minimizing each function defined in ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$, iterating from the stage $\tau=0$ to $\tau=T$. If their minimum is positive at a stage $t$ and there is no intermediate stage $\tau<t$ such that $x_{\tau}$ is outside ${\check{\mathcal{C}}}$, we claim that the initial state $x_{0}$ can safely reach the original target set ${\mathcal{T}}$, under all possible disturbances.
@@
where the first two inequalities are element-wise and the bounds on the dynamics $f$ can be derived using its Lipschitz constant or a Taylor series [^1] For simplicity, we can also use ${\mathcal{X}_{t,\bar{x}_{0}}^{L}}$ to define the convex outer approximation set of the dynamically feasible states in SOCP certifications, i.e., ${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}:={\mathcal{X}_{t,\bar{x}_{0}}^{L}}$. . At a stage $t$, we can verify if $x_{t}\in{\check{\mathcal{T}}}$ by solving a sequence of SOCPs iterating over all $i\in\mathcal{I}_{\check{\mathcal{T}}}$, and checking if their minimum is positive:
${\check{r}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{T}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}$ $P_{i}x_{t}-k_{i}\big{\}}.$ (23)
@@
${\check{c}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{C}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\frac{1}{2}x_{t}^{\top}Q_{i}x_{t}+q_{i}^{\top}x_{t}+b_{i}\big{\}}.\vspace{-0.5em}$ (24)
@@
and $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs, capable of driving all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite disturbances.
Running example (continued). At stage $t$, we evaluate ${\check{r}_{t}^{S}}=\min_{i}\{\check{r}_{t,i}^{S}\}_{i=1}^{6}$ by solving the following SOCPs, where each $\check{r}_{t,i}^{S}$ corresponds a function in the definition of ${\mathcal{T}}$ in (5):
@@
$\check{r}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3-p_{z,t}^{1},$ $\check{r}_{t,6}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3+p_{z,t}^{1}$
@@
$\check{c}_{t,i}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}(-1)^{i}\times p_{z,t}^{1}-p_{y,t}^{1}+0.05,\hskip 5.69046pti\in\{3,4\}.$
We overapproximate the maximum height difference between two drones via $\Delta_{z,t}^{21}:=\max_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}p_{z,t}^{2}-p_{z,t}^{1}$, and consider
$\check{c}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\mathopen{}\mathclose{{}\left\|\begin{bmatrix}p_{x,t}^{1}-p_{x,t}^{2}\\ p_{y,t}^{1}-p_{y,t}^{2}\end{bmatrix}}\right\|_{2}^{2}-(1+\max(\Delta_{z,t}^{21},0))\times 0.2$
Remark 1.
1) SOCP certification can be less conservative than Lipschitz certification when ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$ can represent ${\mathcal{T}}$ and ${\mathcal{C}}$ exactly, and ${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}$ is a subset of the Lipschitz dynamically feasible set ${\mathcal{X}_{t,\bar{x}_{0}}^{L}}$, as defined in Section VI-A, for all $t\leq T$. This is because Lipschitz certification adds extra conservatism when estimating lower bounds of $r(x)$ and $c(x)$ using the Lipschitz constant, as shown in (19) and (20); 2) However, Lipschitz certification is faster to compute than SOCP because calculating ${\check{r}_{t}^{L}}$ and ${\check{c}_{t}^{L}}$ is easier than evaluating ${\check{r}_{t}^{S}}$ and ${\check{c}_{t}^{S}}$; 3) SOCP certification employs convex over-approximations of the dynamically feasible state sets, similar to tube MPC [33]. However, it differs from tube MPC in that we compute the worst-case (nonconvex) constraint violation and target set deviation rather than the control inputs that optimize an objective.
@@
VII Combining reachability learning and certification
@@
Online certification: Let $x$ be an arbitrary state. We can use the RA certificates ${\check{V}^{L}_{\gamma}}(x,T)$ in (21) or ${\check{V}^{S}_{\gamma}}(x,T)$ in (25) as online RA certification methods, verifying if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq{\varepsilon_{x}}\}$ can reach the target set safely despite disturbances. We can compute ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ in real-time (10 Hz in our examples), as shown in Figure 5. It can also be integrated into the safety filter proposed in [10], offering more robust RA certification than [10] by verifying that all states in ${\mathcal{E}}_{x}$ can reach the target set safely, thereby enabling RA capability verification without perfect state estimation.
Offline certification: We consider a finite set of states :=$\mathcal{L}:=\{x^{(i)}\}_{i=1}^{N}$ such that the union of their neighboring sets ${\mathcal{E}}_{x^{(i)}}=\{x:\|x-x^{(i)}\|_{2}\leq{\varepsilon_{x}}\}$ covers the set of states that we aim at certifying. For example, this includes the area near the orange gate in drone racing, as show in Figure 1. We enumerate each element $x\in\mathcal{L}$ and certify whether $\mathcal{E}_{x}\subseteq{\mathcal{R}}$ by checking if ${\check{V}^{L}_{\gamma}}$ in (21) or ${\check{V}^{S}_{\gamma}}$ in (25) is positive. The union ${\mathcal{S}}$ of those certified sets constitutes a subset of the ground truth RA set ${\mathcal{R}}$. For all elements in ${\mathcal{S}}$, we guarantee that they can reach the target set safely under potential disturbances.
@@
Algorithm 1 Certifiable Reachability Learning:
VIII Experiments
@@
We test our reachability learning and certification methods [^2] Experiment code and hardware drone racing video are available at https://
github.com/jamesjingqili/Lipschitz_Continuous_Reachability_Learning.git
@@
in a 12-dimensional drone racing hardware experiment (Figure 3), and a triple-vehicle highway take-over simulation. In the highway simulation, we control one ego vehicle, modeled with nonlinear unicycle dynamics, to safely overtake another vehicle while avoiding a third vehicle driving in the opposite direction, as shown in Figure 4. Figures 1 and 4 demonstrate the high quality of the learned and certified RA sets.
@@
VIII-A Hypothesis 1: Our learned policy has a higher success rate than state-of-the-art constrained RL methods
@@
We compare our learned policy ${\pi_{\theta_{u}}}$ with Deep Deterministic Policy Gradient-Lagrangian (DDPG-L)[35], Soft Actor Critic-Lagrangian (SAC-L) [36], and Constrained Policy Optimization (CPO) [34]. We summarize the results in Table I. The success rate is estimated by computing the ratio of sampled initial states that can reach the target set safely under randomly generated disturbances. Our method achieves a 1.0 success rate when the initial states are sampled from the SOCP certified set, validating the deterministic guarantee that all elements in the certified sets can safely reach the target set despite disturbances.
Initial states uniform sampling Drone racing Highway DDPG-L SAC-L Ours DDPG-L SAC-L CPO Ours In a large bounded state set 0.5716 0.7291 0.7655 0.7512 0.6343 0.5552 0.8782 In the learned reach-avoid set 0.6276 0.8006 0.8889 0.9430 0.9149 0.9566 0.9924 In the SOCP certified set 0.9673 0.9948 1.0000 0.9111 0.8850 0.9451 1.0000
@@
Table I: Success rates table. Our method achieves a $1.0$ success rate when the initial states are sampled from the SOCP certified set. CPO fails to converge for the drone racing experiment due to the complex and nonconvex constraints.
Figure 3: We sampled 50 initial states from the SOCP certified set shown in Figure 1. A few crashes occurred due to insufficient battery charge or Vicon sensor failures caused by natural light. These instances were excluded as outliers. With a fully charged battery and no Vicon system failures, the ego drone successfully overtook the other drone from each of the 50 initial states, despite the latter’s uncertain acceleration. We visualize two hardware experiments in the above subfigures. The remaining 9-dimensional initial state includes $[v_{x,t}^{1},v_{y,t}^{1},v_{z,t}^{1},p_{x,t}^{2},p_{y,t}^{2},p_{z,t}^{2},v_{x,t}^{2},v_{y,t}^{2},v_{z,t}^{2}]=[0,0.7,0,0.4,-2.2,0,0,0.3,0]$.
@@
Figure 4: Highway reachability analysis: In (a), we simulate the nonlinear dynamics with the learned policy ${\pi_{\theta_{u}}}$ and randomly sampled disturbances on other vehicles’ acceleration. The 10-dimensional state space includes $[p_{x,t}^{1},p_{y,t}^{1},v_{t}^{1},\theta_{t}^{1},p_{x,t}^{2},p_{y,t}^{2},v_{y,t}^{2},p_{x,t}^{3},p_{y,t}^{3},v_{y,t}^{3}]$. The $p_{y}$-axis movement of the red and green agents is modeled using double integrator dynamics, while their initial $p_{x}$ positions are sampled randomly and remain stationary during simulation. In (b), we project our learned value function, with $\gamma=0.95$, onto the $(x,y)$ position of the ego vehicle. In (c), we plot the RA set learned using the state-of-the-art method [11, 14] with $\gamma=0.95$. As suggested in [7], annealing $\gamma\to 1$ is necessary for prior works; otherwise, the learned RA sets in prior works are conservative. In (d), we plot our certified RA sets.
@@
VIII-B Hypothesis 2: Our online RA set certification methods can be computed in real-time
Figure 5 shows that ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ can be computed in real-time to certify if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq 0.1\}$ can safely reach the target set, under all potential disturbances. This enables real-time online certification.
Figure 6: We compare the convergence of the critic loss under our new Bellman equation with the baseline from previous works [14, 11], using different $\gamma$ values but identical training parameters. Our critic loss with $\gamma=0.95$ converges faster than with $\gamma=0.9999$, likely due to the Lipschitz continuity of ${V_{\gamma}(x)}$ at $\gamma=0.95$. The training speed is around 1700 steps per second.